Solve (u+1)^3/4=3 where u is a real solution.

Thank you!

Answer by RagingGod(3)

Log both sides of the equation
3/4 LOG (U+1)= LOG 3

LOG (U+1) = 4/3 LOG 3
LOG (U+1) = LOG 3^4/3
U+1 = 3^4/3
U = 3^4/3 - 1
= 3.33 (3 significant figures)